OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 20 Help me please I will give a medal and fan!!

OpenStudy (anonymous):

So far i have this x2 + 2x + y2 + 4y = 20 (x2 + 2x) + (y2 + 4y) = 20

OpenStudy (misty1212):

H!!

OpenStudy (misssunshinexxoxo):

The radius is r = sqrt(25) = 5

OpenStudy (misty1212):

complete mr square twice, once for each term

OpenStudy (anonymous):

I need help figuring it out though, i have no idea whatsoever ho to do this problem.

OpenStudy (anonymous):

@dan815 @Michele_Laino

OpenStudy (anonymous):

Can someone walk me through this please?

OpenStudy (danjs):

for each to complete the square, take the coefficient on the linear x, y term and add to that quantity half of its square

OpenStudy (anonymous):

I honestly have no idea how to do that..

OpenStudy (anonymous):

So do i have to figure out that equation? I honestly dont think ill be able to learn this, ive been trying for about an hour

OpenStudy (danjs):

(x^2 + 2x + 1^2 ) + (y^2 + 4y + 2^2) = 20 + 1^2 + 2^2 you take half the middle term and add the square of it , now you can do this (x+1)^2 + (y+2)^2 = 25

OpenStudy (danjs):

the center of a circle in that form is (x - h)^2 + (y - k)^2 = r^2 here C(h,k) = C(-1 , -2) and r is the square root of 25

OpenStudy (anonymous):

so the radius is 25 and the center of the circle is 5?

OpenStudy (anonymous):

since 5 is the square root of 25

OpenStudy (danjs):

the center is the point (h,k) or ( -1 , -2),, how far each point on the circle is away from that center; the radius, is the square root of 25, or 5

OpenStudy (danjs):

r^2 = 25 r = 5

OpenStudy (anonymous):

i am so confused.....

OpenStudy (danjs):

after you complete the squares, you have an equation in the form of ... (x-h)^2 + (y-k)^2 = r^2 r= radius, and (h,k) = the center point